​​​​​​Question 1. Substitution of given form of solution and hyperbolic functions.Question 1. Substitution of given form of solution and hyperbolic functions. The non-linear ordinary differential equation de
Answer #1

Solution:

we have the equation,

dy dr

we assume y 2

By assuming y-z we can represent the above as a first order equation:

z'=k\sqrt {1 + {z^2}}

\mathrm{d}z=k\sqrt {1 + {z^2}}\mathrm{d}x

dz

\ln \left( {z + \sqrt {1 + {z^2}} } \right)=kx+C

z + \sqrt {1 + {z^2}} =e^{kx+C}

\text{Multiplying both sides of the equation by the conjugent expression } z - \sqrt {1 + {z^2}}

\left( z + \sqrt {1 + {z^2}} \right)\cdot \left( z - \sqrt {1 + {z^2}} \right) = \left( z - \sqrt {1 + {z^2}} \right)e^{kx+C}

z - \sqrt {1 + {z^2}} =-e^{-kx-C}

Adding to the previous equation, we find the expression for z -y iln

2z=e^{kx+C}-e^{-kx-C}

z=\frac{e^{kx+C}-e^{-kx-C}}{2}

z=\sinh \left (kx+C \right )

\frac{\mathrm{d} y}{\mathrm{d} x}=\sinh \left ( kx+C \right )

\text{b)}

\int \mathrm{d} y=\int \sinh \left ( kx+C \right )\mathrm{d} x

cosh (kx C), + C1 y(x)-